# RALAD: Bridging the Real-to-Sim Domain Gap in Autonomous Driving with Retrieval-Augmented Learning

### v1

@@ IV EXPERIMENT
Figure 3: The table displays two sets of nearest neighbor results for KITTI 1, with these results drawn from the CARLA dataset. CARLA 1 stands out as the optimal match for KITTI 1 among the retrieved neighbors. Below the table, a pair of bar graphs illustrate the transmission cost comparison between KITTI 1 and the CARLA features, a lower transmission cost indicates a higher degree of similarity between the images.
To evaluate the effectiveness of the proposed RALAD method, which improves recognition of real-world scenes using simulated data and bridges the gap between simulated and real-world environments, we employed a multi-dataset approach involving KITTI, CARLA datasets, and other real-world images. This method allows for a comprehensive evaluation of the performance of RALAD. The experimental process begins with an overview of the implementation and dataset details, followed by the experimental results of our method, and demonstrate the performance of RALAD retrieval and its applications in BEV perception. Finally, we show the influence of the convex combination ratio on the model’s performance during fine-tuning.
Figure 4: The performance of models with RALAD in CARLA: All models show varying degrees of performance improvement in 3D object detection
@@ IV-A Dataset and Metrics
The KITTI dataset comprises 7481 monocular images from vehicle front cameras, split into 3712 training and 3769 validation images based on Chen et al.’s 3D object detection criteria, with ground truth derived from [5]. For the CARLA dataset, collected via the CARLA 0.9.15 software with robust annotation features, we gathered data from maps like Town01, Town02, and Town07, comprising 500 training, 473 validation, and 227 test images, totaling 1200. Utilizing the RALAD algorithm, we extracted 4066 features from the KITTI and CARLA training sets and fine-tuned the pre-trained Cross View model on KITTI. The 4066 extracted features were allocated to training (2536) and validation (1530) sets. We then assessed the fine-tuned model on both datasets to showcase its real-world and simulated performance, using Mean Intersection over Union (mIOU) and Mean Precision (mAP) as metrics.
Figure 5: Evaluation in the real world. Center: the compared baseline (Cross_view). Right: the baseline enhanced by RALAD (RALAD_Cross_view).
@@ IV-B Experimental Results
We conducted comprehensive experiments using the KITTI and CARLA datasets. To ensure the generalization of RALAD, we selected three models (MonoLayout, Cross View, and DcNet) that are highly regarded in the field of 3D object detection. Detailed experiments were performed on both the original models and their RALAD fine-tuned counterparts in a consistent hardware and software environment.
@@
Metrics Results: Our experimental metrics are shown in Table I, which shows the performance of three models is compared with and without RALAD fine-tuning on KITTI (real-world) and CARLA (simulation) datasets. For MonoLayout, RALAD shows minimal improvement on the KITTI dataset (mIOU from 30.18% to 30.26%, with a slight mAP decrease from 45.91% to 44.98%). However, on the CARLA dataset, it achieves significant improvements (mIOU from 25.26% to 34.13%, mAP from 48.93% to 56.41%). This suggests that RALAD helps more in simulation environments, particularly for this model. For Cross View, the gains on KITTI are modest (mIOU from 38.85% to 39.21%, mAP from 56.64% to 56.49%), while on CARLA, the improvements are much more pronounced (mIOU from 30.55% to 40.82%, mAP from 53.25% to 65.54%). For DcNet, the improvements on KITTI are slight (mIOU from 39.44% to 39.07%, mAP from 58.89% to 57.64%), but the model benefits greatly from RALAD on CARLA (mIOU from 31.09% to 42.11%, mAP from 54.72% to 67.24%). These results suggest that RALAD has a more substantial impact on the simulation dataset, improving both detection accuracy and precision.
Training Overhead: Our RALAD adopts the fine-tuning approach, which can reduce the time cost of re-training existing models, as shown in Table II. To compare the performance of RALAD fine-tuning with standard model re-training, we conducted experiments using 1200 features in a consistent environment. For Cross View, the original model requires 41.54 seconds per epoch on KITTI and 41.33 seconds on CARLA in re-train, while RALAD fine-tuning only needs 5.27 seconds on KITTI and 5.29 seconds on CARLA. In comparison, MonoLayout and DcNet also show significant reductions in training time with RALAD, from 34.84 to 3.81 seconds on KITTI and from 68.07 to 20.23 seconds, respectively. These results highlight RALAD’s efficiency, especially for large-scale adaptation.
Visualizations: To further demonstrate the improvements of RALAD on the CARLA dataset, we re-tested the scene from Figure 1, as shown in Figure 4. There is a significant improvement in detecting vehicles in close-range scenes, especially with DcNet, which successfully identifies vehicles entirely. For distant scenes, there is a qualitative leap in detection, almost matching the ground truth. Given the unique characteristics of real and virtual environments, we reconstructed a nearly identical mapping of the real-world scene in CARLA to minimize external factors. Visualizations were conducted to illustrate the improvements RALAD offers over the original models in the same scenarios. We collected parking lot data using our autonomous vehicle equipped with the open-source Autoware. Universe autonomous driving system for testing. The same positions and parking lot scenarios were recreated in CARLA for comparison. As shown in Figure 5, the fine-tuned models with RALAD provide more accurate vehicle detection in both real and virtual environments when compared to the original BEV perception images. We synthesized the comparison dataset into the video.
IV-C OT Improves The Performance
In this section, we conducted experiments to validate the effectiveness of our OT algorithm in computing the similarity between real-world and virtual-world feature maps. As shown in Figure 3, the algorithm successfully retrieves the most approximate feature map from each dataset. Specifically, a lower OT distance indicates a higher degree of similarity between two feature maps, such as those from KITTI1 and CARLA1 in Group 1 and Group 2. Both groups demonstrate high consistency in key features (e.g., vehicle positions and lane conditions) and achieve the minimum required transmission costs of 62.27×$\times 10^{4}$ and 70.21×$\times 10^{4}$, respectively. These results highlight the capability of our pixel-level OT algorithm to accurately compute the similarity between feature maps in practical applications, enabling effective matching of real and virtual data. The findings are consistent with our proposed theory, confirming the enhanced performance of the improved OT algorithm in similarity calculations, thereby providing a solid foundation for subsequent feature fusion and model training.
IV-D BEV Perception Domain Adaptation
Our RALAD framework is designed to meet the requirements of BEV perception tasks directly and efficiently. When addressing BEV perception, RALAD performs precise feature matching in the target domain, accurately mapping the source and target domains. It retrieves the feature representation most similar to each feature in the source domain, ensuring a high degree of correspondence between the two domains. As shown in Figure 3, this consistency is demonstrated between the KITTI 1 and Carla 1 groups. Subsequently, we leverage these fused feature representations and integrate spatial information from both the source and target domains to fine-tune the final decoder. By adopting a pixel-level OT strategy, this approach achieves cross-domain alignment within the BEV instance space, taking advantage of geometric structures. This strategy significantly improves pixel-level alignment accuracy between the source and target domains, effectively transferring discriminative information from the source domain to the target domain through the feature fusion process. As a result, the model learns more generalized feature representations, enhancing its generalization ability and robustness across different scenarios.
IV-E Convex Combination Ratio
A core aspect of our RALAD framework is the fusion of features extracted from real and virtual data. Different fusion ratios can yield significantly varied results for the model. To explore this, we conducted comparison experiments using four different ratio settings, based on a total of 1800 features. We tested the following combinations of KITTI to CARLA: 0.7:0.3, 0.6:0.4, 0.5:0.5, and 0.4:0.6. As shown in Table III, For the 0.6:0.4 combination, the model achieved a balanced performance, slightly decreasing on KITTI (32.10% mIOU, 54.02% mAP) but showing a substantial improvement on CARLA (35.12% mIOU, 63.33% mAP), making it an optimal trade-off between real and simulated data. In contrast, the 0.7:0.3 combination performed better on KITTI (34.70% mIOU, 54.48% mAP) but much worse on CARLA. As the CARLA ratio increased (0.5:0.5 and 0.4:0.6), performance on CARLA continued to improve, reaching the highest metrics with the 0.4:0.6 combination (36.83% mIOU, 61.39% mAP), but at the cost of reduced performance on KITTI. The experiments indicate that as the ratio of CARLA data increases, the model’s performance on CARLA improves significantly, with the best results seen at the 0.4:0.6 ratio. Conversely, the model’s performance on KITTI decreases as the CARLA ratio increases. The experimental data clearly indicates that 0.6:0.4 is an effective ratio, which not only ensures the performance of the model but also demonstrates good robustness and adaptability. Therefore, in our experiments, we recommend and adopt this ratio for feature fusion in order to achieve more outstanding results in experimental results and application value.
@@ IV-E Convex Combination Ratio
DcNet [7] 39.44 58.89 31.09 54.72
DcNet+RALAD (ours) 39.07 57.64 42.11$\uparrow$ 67.24$\uparrow$
@@
DcNet [7] 68.07 s/epoch 67.97 s/epoch
DcNet+RALAD 20.23 s/epoch 21.08 s/epoch
@@ V CONCLUSIONS
In conclusion, our RALAD model has demonstrated remarkable effectiveness in reducing the gap between real and simulated scenarios. By doing so, it not only maintains the high accuracy achieved in real scenes but also significantly improves the detection accuracy in simulated scenarios. This achievement holds great promise for the field of autonomous driving. Looking ahead, we are determined to conduct further experiments with RALAD in other areas of autonomous driving.

### v2

@@ IV EXPERIMENT
To evaluate the effectiveness of RALAD , we conduct experiments on three models trained on the KITTI dataset, a widely recognized benchmark in autonomous driving. We first present the metrics and dataset details, followed by experimental results showcasing the improvements achieved by RALAD, then we analyze the training efficiency of RALAD and provide a qualitative visualization on one of the models. Finally, we conduct a series of ablation studies to further validate our approach.
@@ IV-A Dataset and Metrics
The KITTI dataset comprises 7481 monocular images from vehicle front cameras, split into 3712 training and 3769 validation images based on [20], with ground truth derived from [5]. For the CARLA dataset, collected via the CARLA 0.9.15 software with robust annotation features, we gathered data from maps like Town01, Town02, and Town07, comprising 500 training, 473 validation, and 227 test images, totaling 1200. Utilizing the RALAD, we extracted 4066 features from the KITTI and CARLA training sets and fine-tuned the models on KITTI, the features were allocated to training (2536) and validation (1530) sets. We then assessed the fine-tuned models on both datasets to showcase its real-world and simulated performance, using Mean Intersection over Union (mIOU) and Mean Precision (mAP) as metrics.
@@ IV-B Experimental Results
Our experimental metrics are presented in Table I, demonstrating the effectiveness of RALAD across all three models. RALAD not only preserves the models’ accuracy on real-world data but also significantly improves their performance on simulated data, achieving recognition capabilities comparable to or even exceeding those on real data. Furthermore, we observe that RALAD provides greater performance gains for newer models, highlighting its effectiveness in enhancing state-of-the-art architectures.
@@
MonoLayout: RALAD shows minimal improvement on the KITTI dataset (mIOU from 30.18% to 30.26%, with a slight mAP decrease from 45.91% to 44.98%). However, on the CARLA dataset, it achieves significant improvements (mIOU from 25.26% to 34.13%, mAP from 48.93% to 56.41%). This suggests that RALAD helps more in simulation environments, particularly for this model.
CrossView: the gains on KITTI are modest (mIOU from 38.85% to 39.21%, mAP from 56.64% to 56.49%), while on CARLA, the improvements are much more pronounced (mIOU from 30.55% to 40.82%, mAP from 53.25% to 65.54%).
DctNet: the improvements on KITTI are slight (mIOU from 39.44% to 39.07%, mAP from 58.89% to 57.64%), but the model benefits greatly from RALAD on CARLA (mIOU from 31.09% to 42.11%, mAP from 54.72% to 67.24%). These results suggest that RALAD has a more substantial impact on the simulation dataset, improving both detection accuracy and precision.
@@
DctNet + $l_{2}$ 39.05 60.69 37.89 61.15
DctNet [7] 39.44 58.89 31.09 54.72
DctNet+RALAD (ours) 39.07 57.64 42.11$\uparrow$ 67.24$\uparrow$
@@
DctNet [7] 68.07 s/epoch 67.97 s/epoch
DctNet+RALAD 20.23 s/epoch 21.08 s/epoch
IV-C Training Overhead
Our RALAD adopts the fine-tuning approach, which can reduce the time cost of re-training existing models, as shown in Table II. To compare the performance of RALAD fine-tuning with standard model re-training, we conducted experiments using 1200 features in a consistent environment. For CrossView, the original model requires 41.54 seconds per epoch on KITTI and 41.33 seconds on CARLA in re-train, while RALAD fine-tuning only needs 5.27 seconds on KITTI and 5.29 seconds on CARLA. In comparison, MonoLayout and DctNet also show significant reductions in training time with RALAD, from 34.84 to 3.81 seconds on KITTI and from 68.07 to 20.23 seconds, respectively. These results highlight RALAD’s efficiency, especially for large-scale adaptation.
IV-D Visualizations
Visualization of Retrival Results: As shown in Figure 5, we visualized the nearest sets of feature maps retrieved by RALAD, to verify that our method can retrieve highly approximate simulated and real-world feature maps. It can be seen that the method successfully retrieves the most approximate feature map from each dataset. Specifically, a lower OT distance indicates a higher degree of similarity between two feature maps, such as those from KITTI1 and CARLA1. Both groups demonstrate high consistency in key features (e.g., vehicle positions and lane conditions) and have the minimum transmission costs of 62.27×$\times 10^{4}$ and 70.21×$\times 10^{4}$, respectively. These results highlight the capability of RALAD to accurately compute the similarity between feature maps in practical applications, enabling effective matching of real and virtual data.
Figure 6: The performance of models with RALAD in simulated rainy weather: All models show varying degrees of performance improvement under simulated rainy weather conditions.
Figure 7: Evaluation in the real world: Center: the compared baseline (Cross_view). Right: the baseline enhanced by RALAD (RALAD_Cross_view).
Simulation and Real Autonomous Vehicle Experiments: To further demonstrate the improvements of RALAD on CARLA dataset, we re-tested the scene from Figure 1, as shown in Figure 6, there is a significant improvement in detecting vehicles in simulation, especially with DctNet, which successfully identifies vehicles entirely. In addition, we collected parking lot data for testing using an autonomous vehicle equipped with the open-source Autoware.Universe autonomous driving system. Considering the unique characteristics of the real and simulated environments, we reconstructed a scene in CARLA that closely resembles the real world to minimize external factors, thereby demonstrating the improvements provided by RALAD compared to the original model. As shown in Figure 7, the models with RALAD provide more accurate detection in both real and simulated environments compared to the original performances.
@@ IV-D Visualizations
IV-E Ablation Experiment
OT Ablation: In this section, we conducted an ablation study on the OT algorithm by separately preserving the $l_{2}$ distance for experiments. The results are shown in Table I. It can be seen that using the $l_{2}$ distance to establish a mapping relationship does not show a significant difference compared to the baseline on KITTI, and also demonstrates a improvement on CARLA. However, when compared to using RALAD, the gap is extremely noticeable. The CrossView using $l_{2}$, as opposed to the OT, has a slight decrease on KITTI (mIOU from 39.21% to 37.29%, mAP from 56.49% to 54.29%), and a significant decrease on CARLA (mIOU from 40.82% to 34.28%, mAP from 65.54% to 53.27%). And DctNet using $l_{2}$ also shows a significant decrease on the CARLA dataset (mIOU from 42.11% to 37.89%, mAP from 67.24% to 61.15%), as opposed to the OT. This experiments clearly indicates that the OT algorithm is more accurate than the $l_{2}$ algorithm in establishing a mapping relationship between real and simulated environments, enabling the model to better bridge the gap between reality and simulation.
Convex Combination Ratio: In the feature merging stage of RALAD, different fusion ratios can produce varied results for the model. To explore this, we conducted comparison experiments with four different ratio settings, validating them on the KITTI and CARLA datasets, using a total of 1800 features. As shown in Table III, the 0.6:0.4 combination achieved a balanced performance, slightly decreasing on KITTI, but showing a substantial improvement on CARLA, making it an optimal trade-off between real and simulated data. In contrast, the 0.7:0.3 combination performed better on KITTI , but much worse on CARLA. As the CARLA ratio increased, performance on CARLA continued to improve, reaching the highest metrics with the 0.4:0.6 combination, but at the cost of reduced performance on KITTI, so we recommend and adopt this ratio for feature fusion to obtain better results.
IV-F Experiments Conclusion
  • As shown in Table I and II, RALAD demonstrated exceptional capability in bridging the real and simulated gap. Ensuring that the learned knowledge is not forgotten, the performance of all three models in simulation improved by approximately 10%. Compared to using the $l_{2}$ distance, RALAD performed better, with a relative improvement about 5% in simulation. In addition, RALAD also achieved a significant reduction in training time.
  • In section IV-D, we clearly observed that the model’s performance significantly improved in simulation, after using RALAD. To further validate its effectiveness, we also used real autonomous driving car and conducted tests in parking lots. The results showed that the model’s performance in the real world also showed a remarkable boost, fully demonstrating the outstanding effectiveness of RALAD.
@@ V CONCLUSIONS
RALAD demonstrates a significant success in bridging the gap between real and simulated driving scenarios. By introducing RALAD, the model effectively constructs a mapping between the real and simulated environments. Experimental results show that RALAD not only maintains high accuracy in reality but also substantially improves detection performance in simulation. Additionally, the fine-tuning strategy employed by the model significantly enhances training efficiency. Overall, RALAD offers a more efficient and reliable solution for autonomous driving testing, advancing the development of autonomous driving technology. Future research could explore applying RALAD to a broader range of tasks and scenarios in autonomous driving.